# Topologically noise robust network steering without inputs

### v1

@@ Abstract
Introduction— Quantum non-locality is one of the most remarkable features of quantum mechanics that defies our classical intuitions about the world. It refers to the ability of separated quantum systems to set-up correlations that are incompatible with assumptions of local causality. This quantum property was first conceptualized in the celebrated work of Einstein, Podolsky and Rosen [1], following which Bell in 1964 [2, 3] formulated a theoretical test to observe non-classical correlations. It was then experimentally verified [4, 5, 6, 7] and is now recognized as a fundamental aspect of quantum mechanics. Quantum non-locality has now found applications in the study of a wide variety of topics such as cryptography, quantum teleportation, quantum communication, and quantum computing (refer to [8] for a review). Quantum non-locality can be re-formulated, with an additional assumption, into quantum steering, where an observer is able to remotely influence a distant quantum system. Quantum steering was initially introduced by Schrodinger [9] and was rigorously formulated in [10]. The additional assumption is that one of the parties is trusted to perform fixed quantum measurements.
@@
The notion of quantum steering in networks was introduced in [16] where some nodes in the network were assumed to be trusted in the sense that local subsystems at these nodes could be precisely known, or equivalently, the parties at the trusted nodes could perform a tomographically complete set of measurements on their incoming subsystems. The minimal scenario where one could observe quantum steering in networks was introduced in [17] and is known as swap-steering. The scenario comprised of two sources and two parties without inputs, where one of the parties is trusted to perform a single bell measurement. Unlike the case in [15], to demonstrate swap-steering, a linear witness was constructed in [17]. This makes swap-steering much easier to observe in experiments. Moreover, it was also recently demonstrated that every entangled state [18] and every entangled measurement [19] can be used to observe swap-steering.
@@
I Triangle network swap-steering
Let us begin by describing the triangle network steering scenario. Each party (or node) in the network receives a subsystem from each source connected to it on which the party performs a single four-outcome measurement. The outcomes are denoted as $a_{n}$ for the $n^{th}$ node such that $n=0,1,2,3$. One of the measurement parties is trusted here implying that the measurement performed on it’s subsystem is known (see Fig. 1). Here we trust that $A_{1}$ performs the measurement corresponding to the Bell basis given by $N_{A_{1}}=\{|\phi_{+}\rangle\!\!\langle\phi_{+}|,|\phi_{-}\rangle\!\!\langle\phi_{-}|,|\psi_{+}\rangle\!\!\langle\psi_{+}|,|\psi_{-}\rangle\!\!\langle\psi_{-}|\}$ where
|ϕ±ketsubscriptitalic-$|\phi_{\pm}\rangle$ $=$ $\frac{1}{\sqrt{2}}\left(|00\rangle\pm|11\rangle\right)$
@@ I Triangle network swap-steering
Here $A_{i}^{j}$ denotes the $j^{th}$ subsystem received by the $i^{th}$ node and $N_{a_{1}}^{A_{1}}/M_{a_{i}}^{A_{i}}$ denotes trusted/untrusted measurement elements of $A_{i}$ corresponding to the outcome $a_{i}$, are positive and $\sum_{a_{i}}M_{A_{i}}^{a_{i}}=\sum_{a_{1}}N_{A_{1}}^{a_{1}}=\mathbbm{1}$. It is important to recall that the measurement parties can not communicate during the experiment.
I.1 SOHS models
The notion of unsteerability in the simplest quantum network was introduced in [17] and is based on two assumptions, namely outcome-independence and separable quantum sources. Any hidden state model satisfying the above two assumptions is termed separable outcome-independent hidden state (SOHS) model. Let us here modify the assumptions of [17] for the triangle network with one trusted node. For this purpose, let us begin by considering the most general model to reproduce any correlation $p\left(\{a\}_{3}\right)$ in the triangle network [see Fig. 1] given by
$p\left(\{a\}_{3}\right)=\sum_{\lambda_{1},\lambda_{2},\lambda_{3}}p(a_{1},a_{2},a_{3}|\lambda_{1},\lambda_{2},\lambda_{3})p(\lambda_{1},\lambda_{2},\lambda_{3})$ (3)
@@ I.1 SOHS models
$\sum_{\lambda_{1},\lambda_{2},\lambda_{3}}p(a_{1}|\lambda_{1},\lambda_{3})p(a_{2},a_{3}|\lambda_{1},\lambda_{2},\lambda_{3},a_{1})p(\lambda_{1},\lambda_{2},\lambda_{3}).$ (4)
@@
Assumption 2 (Separable sources).
@@
If correlations in the triangle network do not admit a SOHS model, then the correlations are swap-steerable. Now, from the expression (13), it is clear that if the local state $\rho_{\lambda_{1},\lambda_{n}}$ is not separable, then such correlations do not admit a SOHS model. This is possible if the joint state prepared by the sources $S_{1},S_{n}$ already sends some entangled state to the trusted party and thus the assumption of separable sources is violated. However, based on the current understanding of the physical world, it should always be possible to prepare two sources that do not produce entangled joint systems. Considering that the sources are separable, this implies that outcome-independence is violated which is detected by the trusted party as his local state is entangled due to the measurement of other parties. Thus, other parties could steer or change the trusted party’s local state from separable to entangled using local operations on spatially separated labs.
@@ I.2 Swap-Steering
[…]
Recall here that Alice is trusted and performs the Bell-basis measurement with elements given in (I). Let us now find the maximum value that can be achieved using correlations that admit a SOHS model (11).
@@
The maximum value $\beta_{SOHS}$ of the three party swap-steering functional $W_{3}$ (I.2) that can be obtained using a SOHS model is $\frac{1}{2}$.
The proof of the above is stated in Fact 1 of the Appendix. Now, consider the following quantum experiment. Each source prepares the state $|\psi_{i}\rangle=|\phi_{+}\rangle_{A_{i}^{2}A_{i+1}^{1}}$ and each node measures in the same basis as the trusted node, that is, $M_{A_{i}}=\{|\phi_{+}\rangle\!\!\langle\phi_{+}|,|\phi_{-}\rangle\!\!\langle\phi_{-}|,|\psi_{+}\rangle\!\!\langle\psi_{+}|,|\psi_{-}\rangle\!\!\langle\psi_{-}|\}_{A_{i}^{1}A_{i}^{2}}$ where the corresponding states are given in (I). Using these states and measurements, one can simply evaluate the steering functional $W_{3}$ (I.2) to get the value $1$. Notice that this is also the algebraic bound of $W_{3}$. Thus, we conclude that the experiment described generates the quantum correlation that maximally violates the three-party swap-steering inequality. It is also interesting to observe here that with the given quantum states and measurements, the post-measured state at the trusted node $A_{1}$ is one of the four bell states (I). Consequently, the state with the trusted party is steered from a separable state to an entangled one due to the measurement of the other parties and is detected via the swap-steering witness.
I.3 Necessary conditions for swap-steering
@@
If any one of the states generated by the sources or the measurement of any untrusted node is separable, then the maximum value that can be obtained of the swap-steering functional $W_{3}$ (I.2) is $\frac{1}{2}$.
The proof of this fact is added in the Appendix. The key insight here is that in the experiment described above for the maximal violation of the SOHS bound, the sources set up a chain of entanglement between the nodes which perform entangled measurements to swap the entanglement between the pairs of systems, resulting in an entangled state at the trusted node post measurement at all the other nodes. However, when any of the states or measurements are separable, there is a break in the link, thus, entanglement cannot be swapped across the network to the trusted node’s system. Thus, steering functional is bounded by $\frac{1}{2}$.
I.4 Weak self-testing in the triangle network
@@
Consider again the triangle scenario described above in Fig. 1. Unlike the self-testing result presented in [17] here we assume that the sources are identical, that is, the sources generate the same state. Let us also remark here that we consider that the states generated by the sources are pure and the measurements are projective. In most of the self-testing statements in literature based on standard Bell or steering scenarios, this assumption holds quite well, however, there might be some exceptions as noted in [47]. Consequently, the certification states and the measurements are weaker compared to the standard self-testing protocols. Let us now state our self-testing result in the triangle network scenario with one trusted party.
@@
Assume that Alice, Bob and Charlie observe that the correlations $p(a,b,c)$ in the steering functional $W_{3}$ (I.2) are equal along with $W_{3}=1$ such that trusted Alice performs the Bell-basis. Then, the following statements hold true:
@@ 1.    There exist unitary transformation $U$, such that the state generated by the sources are certified as
$(\mathbbm{1}_{A_{2}^{1}}\otimes U_{A^{2}_{3}})|\psi_{A_{2}^{1}A_{3}^{2}}\rangle=|\phi^{+}\rangle.$ (16)
@@ 2.    The measurement of Bob $\{M_{A_{2}}^{a_{2}}\}$ and Charlie $\{M_{A_{3}}^{a_{3}}\}$ is certified as
The proof of the above fact is given in the Appendix.
II Extensions
II.1 $n$-ring scenario
Let us now consider a generalisation of the triangle network [Fig. 1] to a network with arbitrary number of nodes Fig. 2. As previously done for the triangle, we trust node $A_{1}$ to measure in the Bell basis (I). Generalising the notion of swap-steerability to this network, any correlations $p(\{a\}_{n})$ on $n-$ring network (ring network with $n$ nodes) that admits a SOHS model must be of the form
@@ II.1 $n$-ring scenario
Let us now construct a witness to observe swap-steering in the $n-$ring network. For this purpose, consider the network with $n-1$ nodes with the corresponding witness given by $W_{n-1}=\sum_{\{a\}_{n-1}}c_{\{a\}_{n-1}}p(a_{1},\ldots,a_{n-1})$. Then, to obtain the witness $W_{n}$ corresponding to the network with $n$ nodes, in the witness $W_{n-1}$ the outcome of the $n-1$ party $a_{n-1}$ is "reverse coarse-grained" or refined as
00 0 $\rightarrow(0,0),\quad(1,1),\quad(2,2),\quad(3,3)$
11 11 $\rightarrow(0,1),\quad(1,0),\quad(2,3),\quad(3,2)$
22 22 $\rightarrow(0,2),\quad(2,0),\quad(1,3),\quad(3,1)$
33 33 $\rightarrow(0,3),\quad(3,0),\quad(1,2),\quad(2,1).$ (19)
Notice that if we consider the witness $W_{2}$ in [17] in the network with two nodes $n=2$, the witness $W_{3}$ (I.2) can be obtained from it using the above "reverse" coarse-graining (II.1). Following the above procedure, the witness in the network with $n$ nodes is given by $W_{n}=\sum_{\{a\}_{n}}c_{\{a\}_{n}}p(\{a\}_{n})$ where
[…]
@@
The proof of the above is stated in Fact 1 of the appendix where we find the upper bound for more general correlations admitting a topologically robust SOHS model stated below (II.2). Consider now the following quantum experiment. Each source prepares the state $|\psi_{i}\rangle=|\phi_{+}\rangle_{A_{i}^{2}A_{i+1}^{1}}$ and each node measures in the same basis as the trusted node, that is, $M_{A_{n}}=\{|\phi_{+}\rangle\!\!\langle\phi_{+}|,|\phi_{-}\rangle\!\!\langle\phi_{-}|,|\psi_{+}\rangle\!\!\langle\psi_{+}|,|\psi_{-}\rangle\!\!\langle\psi_{-}|\}_{A_{n}^{1},A_{n}^{2}}$ where the corresponding states are given in (I). By assuming that $W_{n-1}=1$ using the above quantum strategy, one can prove by induction that $W_{n}=1$ using the same strategy. This is understood as the Bell basis measurement at the n-th node simply swapping entanglement to the (n-1)-th and 1-st node in a manner consistent with the coarse-graining scheme (II.1). These states will in fact correspond to a Bell basis and thus it reduces to a network with $n-1$ nodes with the above-stated quantum states and measurements. Notice that this is also the algebraic bound of $W_{n}$. Thus, the experiment described above is an example of the maximal quantum violation of the steering inequality. Similar to the triangle network, all the sources must necessarily distribute entangled states and all the nodes must measure an entangled basis to generate $n-$ring swap-steering correlations.
II.2 Topologically Robust Swap Steering
Let us now consider the scenario where the $n-1$ untrusted parties are conncected via addtional sources in an arbitrary way. The strongest such scenario would be one where the untrusted parties and the sources connecting them are morphed into one single party that receives the hidden variables $\lambda_{1}$ and $\lambda_{n}$, and outputs a distribution of $n-1$ outcomes. Here we will prove that the SOHS bound for this scenario will remain the same. Note that we still require that the trusted party receives a separable two-qubit state. This will also serve as the bound for scenarios in which the untrusted parties posses weaker connections among themselves, for example the ring scenario considered previously. The SOHS models for this scenario will be termed as topologically robust swap-steering (TSOHS) models and will be given by

### v2

@@ Abstract
I Introduction
Quantum non-locality is one of the most remarkable features of quantum mechanics that defies our classical intuitions about the world. It refers to the ability of separated quantum systems to set-up correlations that are incompatible with assumptions of local causality. This quantum property was first conceptualized in the celebrated work of Einstein, Podolsky and Rosen [1], following which Bell in 1964 [2, 3] formulated a theoretical test to observe non-classical correlations. It was then experimentally verified [4, 5, 6, 7] and is now recognized as a fundamental aspect of quantum mechanics. Quantum non-locality has now found applications in the study of a wide variety of topics such as cryptography, quantum teleportation, quantum communication, and quantum computing (refer to [8] for a review). Quantum non-locality can be re-formulated, with an additional assumption, into quantum steering, where an observer is able to remotely influence a distant quantum system. Quantum steering was initially introduced by Schrodinger [9] and was rigorously formulated in [10]. The additional assumption is that one of the parties is trusted to perform fixed quantum measurements.
@@ I Introduction
The notion of quantum steering in networks was introduced in [16] where some nodes in the network were assumed to be trusted in the sense that local subsystems at these nodes could be precisely known, or equivalently, the parties at the trusted nodes could perform a tomographically complete set of measurements on their incoming subsystems. The minimal scenario where one could observe quantum steering in networks without inputs was introduced in [17] and is known as swap-steering. The scenario comprised of two sources and two parties without inputs, where one of the parties is trusted to perform a single bell measurement. Unlike the case in [15], to demonstrate swap-steering, a linear witness was constructed in [17]. This makes swap-steering much easier to observe in experiments. Moreover, it was also recently demonstrated that every entangled state [18] and every entangled measurement [19] can be used to observe swap-steering.
@@
II Triangle network swap-steering
Let us begin by describing the triangle network steering scenario. Each party (or node) in the network receives a subsystem from each source connected to it on which the party performs a single four-outcome measurement. The outcomes are denoted as $a_{n}$ for the $n^{th}$ node such that $a_{n}=0,1,2,3$ and $n=1,2,3$. One of the measurement parties is trusted here implying that the measurement performed on it’s subsystem is known (see Fig. 1). Here we trust that $A_{1}$ performs the measurement corresponding to the Bell basis given by $N_{A_{1}}=\{|\phi_{+}\rangle\!\!\langle\phi_{+}|,|\phi_{-}\rangle\!\!\langle\phi_{-}|,|\psi_{+}\rangle\!\!\langle\psi_{+}|,|\psi_{-}\rangle\!\!\langle\psi_{-}|\}$ where
$|\phi_{\pm}\rangle$ $=$ $\frac{1}{\sqrt{2}}\left(|00\rangle\pm|11\rangle\right)$
@@ II Triangle network swap-steering
Here $A_{i}^{j}$ denotes the $j^{th}$ subsystem received by the $i^{th}$ node and $N_{A_{1}}^{a_{1}}/M^{a_{i}}_{A_{i}}$ denotes trusted/untrusted measurement elements of $A_{i}$ corresponding to the outcome $a_{i}$, are positive and $\sum_{a_{i}}M_{A_{i}}^{a_{i}}=\sum_{a_{1}}N_{A_{1}}^{a_{1}}=\mathbbm{1}$. It is important to recall that the measurement parties can not communicate during the experiment.
SOHS models— The notion of unsteerability in the simplest quantum network was introduced in [17] and is based on two assumptions, namely outcome-independence and separable quantum sources. Any hidden state model satisfying the above two assumptions is termed separable outcome-independent hidden state (SOHS) model. Let us here modify the assumptions of [17] for the triangle network with one trusted node. For this purpose, let us begin by considering the most general model to reproduce any correlation $p\left(\{a\}_{3}\right)$ in the triangle network [see Fig. 1] given by
$p\left(\{a\}_{3}\right)=\sum_{\lambda_{1},\lambda_{2},\lambda_{3}}p(a_{1},a_{2},a_{3}|\lambda_{1},\lambda_{2},\lambda_{3})p(\lambda_{1},\lambda_{2},\lambda_{3})\ \$(3)
@@
$\sum_{\lambda_{1},\lambda_{2},\lambda_{3}}p(a_{1}|\lambda_{1},\lambda_{3})p(a_{2},a_{3}|\lambda_{1},\lambda_{2},\lambda_{3},a_{1})p(\lambda_{1},\lambda_{2},\lambda_{3}).\ \$(4)
@@
Assumption 2 (Separate sources).
@@
If correlations in the triangle network do not admit a SOHS model, then the correlations are swap-steerable. Now, from the expression (13), it is clear that if the local state $\rho_{\lambda_{1},\lambda_{3}}$ is not separable, then such correlations do not admit a SOHS model. This is possible if the joint state prepared by the sources $S_{1},S_{3}$ already sends some entangled state to the trusted party and thus the assumption of separable sources is violated. However, based on the current understanding of the physical world, it should always be possible to prepare two sources that do not produce entangled joint systems. Considering that the sources are separable, this implies that outcome-independence is violated which is detected by the trusted party as his local state is entangled due to the measurement of other parties. Thus, other parties could steer or change the trusted party’s local state from separable to entangled using local operations on spatially separated labs.
@@ II.1 Swap-Steering
$c_{\{a\}_{3}}=\begin{cases}1\ \mathrm{if}\ a_{1}-a_{2}-(-1)^{a_{2}}a_{3}\ \mathrm{mod}\ 4=0\\ 0\ \mathrm{otherwise}.\end{cases}$ (15)
Recall here that Alice is trusted and performs the Bell-basis measurement with elements given in (II). Let us now find the maximum value that can be achieved using correlations that admit a SOHS model (11).
@@
The maximum value $\beta_{SOHS}$ of the three party swap-steering functional $W_{3}$ (II.1) that can be obtained using a SOHS model is $\frac{1}{2}$.
The proof of the above is stated in Fact 1 of the Appendix. Now, consider the following quantum experiment. Each source prepares the state $|\psi_{i}\rangle=|\phi_{+}\rangle_{A_{i}^{2}A_{i+1}^{1}}$ and each node measures in the same basis as the trusted node, that is, $M_{A_{i}}=\{|\phi_{+}\rangle\!\!\langle\phi_{+}|,|\phi_{-}\rangle\!\!\langle\phi_{-}|,|\psi_{+}\rangle\!\!\langle\psi_{+}|,|\psi_{-}\rangle\!\!\langle\psi_{-}|\}_{A_{i}^{1}A_{i}^{2}}$ where the corresponding states are given in (II). Using these states and measurements, one can simply evaluate the steering functional $W_{3}$ (II.1) to get the value $1$. Notice that this is also the algebraic bound of $W_{3}$. Thus, we conclude that the experiment described generates the quantum correlation that maximally violates the three-party swap-steering inequality. It is also interesting to observe here that with the given quantum states and measurements, the post-measured state at the trusted node $A_{1}$ is one of the four bell states (II). Consequently, the state with the trusted party is steered from a separable state to an entangled one due to the measurement of the other parties and is detected via the swap-steering witness.
II.2 Necessary conditions for swap-steering
@@
If any one of the states generated by the sources or the measurement of any untrusted node is separable, then the maximum value that can be obtained of the swap-steering functional $W_{3}$ (II.1) is $\frac{1}{2}$.
The proof of the above statement is stated in Fact 3 of the Appendix. The key insight here is that in the experiment described above for the maximal violation of the SOHS bound, the sources set up a chain of entanglement between the nodes which perform entangled measurements to swap the entanglement between the pairs of systems, resulting in an entangled state at the trusted node post measurement at all the other nodes. However, when any of the states or measurements are separable, there is a break in the link, thus, entanglement cannot be swapped across the network to the trusted node’s system. Thus, steering functional is bounded by $\frac{1}{2}$.
II.3 Weak self-testing in the triangle network
@@
Consider again the triangle scenario described above in Fig. 1. Unlike the self-testing result presented in [17] here we assume that the sources are identical, that is, the sources generate the same state. Let us also remark here that we consider that the states generated by the sources are pure and the measurements are projective. This is a standard assumption is most Bell or steering based self-testing, however, there might be some exceptions as noted in [47]. Consequently, the certification of the states and measurements are weaker compared to the standard self-testing protocols. Let us now state our self-testing result in the triangle network scenario with one trusted party.
@@
Assume that Alice, Bob and Charlie observe that the correlations $p(a,b,c)$ in the steering functional $W_{3}$ (II.1) are equal along with $W_{3}=1$ such that trusted Alice performs the Bell-basis. Considering identical sources, such that the states generated by them are pure and measurements of untrusted parties are projective, the following statements hold true:
@@ 1. There exist unitary transformation $U$, such that the state generated by the sources are certified as
$(\mathbbm{1}_{A_{2}^{1}}\otimes U_{A^{2}_{3}})|\psi_{A_{2}^{1}A_{3}^{2}}\rangle=|\phi^{+}\rangle.\ \$(16)
@@ 2. The measurement of Bob $\{M_{A_{2}}^{a_{2}}\}$ and Charlie $\{M_{A_{3}}^{a_{3}}\}$ is certified as
The proof of the above fact is given in the Appendix. Let us now consider a generalisation of the triangle network [Fig. 1] to a network with arbitrary number of nodes Fig. 2.
III $n$-ring scenario
As previously done for the triangle, we trust node $A_{1}$ to measure in the Bell basis (II). Generalising the notion of swap-steerability to this network, any correlations $p(\{a\}_{n})$ on $n$-ring network (ring network with $n$ nodes) that admits a SOHS model must be of the form
@@ III $n$-ring scenario
Let us now construct a witness to observe swap-steering in the $n$-ring network. For this purpose, consider the network with n1n-1 nodes with the corresponding witness given by $W_{n-1}=\sum_{\{a\}_{n-1}}c_{\{a\}_{n-1}}p(a_{1},\ldots,a_{n-1})$. Then, to obtain the witness $W_{n}$ corresponding to the network with $n$ nodes, in the witness $W_{n-1}$ the outcome of the n1n-1 party $a_{n-1}$ is "reverse coarse-grained" or refined as
$0$ $\rightarrow(0,0),\quad(1,1),\quad(2,2),\quad(3,3)$
$1$ $\rightarrow(0,1),\quad(1,0),\quad(2,3),\quad(3,2)$
$2$ $\rightarrow(0,2),\quad(2,0),\quad(1,3),\quad(3,1)$
$3$ $\rightarrow(0,3),\quad(3,0),\quad(1,2),\quad(2,1).$ (19)
Notice that if we consider the witness $W_{2}$ in [17] in the network with two nodes $n=2$, the witness $W_{3}$ (II.1) can be obtained from it using the above "reverse" coarse-graining (III). Following the above procedure, the witness in the network with $n$ nodes is given by $W_{n}=\sum_{\{a\}_{n}}c_{\{a\}_{n}}p(\{a\}_{n})$ where
$c_{\{a\}_{n}}=\begin{cases}1\ \mathrm{if}\ a_{1}-a_{2}-\sum_{i=3}^{n}(-1)^{\sum_{k=2}^{i-1}a_{k}}a_{i}\ \mathrm{mod}\ 4=0\\ 0\ \mathrm{otherwise}.\end{cases}$ (20)
@@
The proof of the above is stated in Fact 1 of the appendix where we find the upper bound for more general correlations admitting a topologically robust SOHS model stated below (III.1). Consider now the following quantum experiment. Each source prepares the state $|\psi_{i}\rangle=|\phi_{+}\rangle_{A_{i}^{2}A_{i+1}^{1}}$ and each node measures in the same basis as the trusted node, that is, $M_{A_{n}}=\{|\phi_{+}\rangle\!\!\langle\phi_{+}|,|\phi_{-}\rangle\!\!\langle\phi_{-}|,|\psi_{+}\rangle\!\!\langle\psi_{+}|,|\psi_{-}\rangle\!\!\langle\psi_{-}|\}_{A_{n}^{1},A_{n}^{2}}$ where the corresponding states are given in (II). By assuming that $W_{n-1}=1$ using the above quantum strategy, one can prove by induction that $W_{n}=1$ using the same strategy. This is understood as the Bell basis measurement at the $n$-th node simply swapping entanglement to the (n1)(n-1)-th and 1-st node in a manner consistent with the coarse-graining scheme (III). These states will in fact correspond to a Bell basis and thus it reduces to a network with n1n-1 nodes with the above-stated quantum states and measurements. This is also proven in Fact 2 of the appendix. Notice that this is also the algebraic bound of $W_{n}$. Thus, the experiment described above is an example of the maximal quantum violation of the steering inequality. Similar to the triangle network, all the sources must necessarily distribute entangled states and all the nodes must measure an entangled basis to generate $n$-ring swap-steering correlations.
III.1 Topologically Robust Swap Steering
Let us now consider the scenario where the n1n-1 untrusted parties are conncected via addtional sources in an arbitrary way. The strongest such scenario would be one where the untrusted parties and the sources connecting them are morphed into one single party that receives the hidden variables $\lambda_{1}$ and $\lambda_{n}$, and outputs a distribution of n1n-1 outcomes [see Fig. 2]. Here we will prove that the SOHS bound for this scenario will remain the same. Note that we still require that the trusted party receives a separable two-qubit state. This will also serve as the bound for scenarios in which the untrusted parties posses weaker connections among themselves, for example the ring scenario considered previously. The SOHS models for this scenario will be termed as topologically robust SOHS (TSOHS) models and will be given by